11. The following shows the age distribution in years of 40 youths at a youth rally.


Age in years (x)

Frequency (f)

7 − 12. 4


12 − 17


17 − 22. 15


22 − 27


27 − 32. 5


Total. 40

Given that the median age is 19.8 years:

i.

Find the missing frequencies to the nearest whole number.

ii.

Find the mode of this distribution.

iii.

How many youths are older than 17 years?

iv.

Include a column for relative frequencies.

v.

Hence, or otherwise, construct a pie chart for this data.


1
Expert's answer
2021-05-07T10:01:57-0400

Given, Median=19.8


(i) Median "M=l+\dfrac{(\frac{N}{2}-cf)}{f}h"

"\Rightarrow 19.8=17+\dfrac{\frac{40}{2}-(4+f_1)}{15}\times 5"


"\Rightarrow 2.8=\dfrac{16-f_1}{3}\Rightarrow f_1=16-8.4=7.6"


Also, "24+f_1+f_2=40"

"f_2=40-24-7.6=8.4"


Missing frequencies are 7.6 and 8.4


(ii) Mode "=l+\dfrac{(f_m-f_{m-1})}{2f_m-f_{m-1}-f_{m+1}}h"

"=17+\dfrac{15-7.6}{2\times 15-7.6-8.4}\times 5"


"=17+2.8=19.8"


(iii) Number of youths older than 17 years are "= 15+f_2+5=20+8.4=28.4"


(iv) Relative frequency table is included in the above chart.


(v) Pie chart for the given distribution is-